# Exploring specialization and sensitivity of convolutional neural networks in the context of simultaneous image augmentations

### v1

@@ 3.3 Cutting off submodules of a network
Corresponding unitwise-averaged correlation matrices are presented in Figures S7.7(a)–(d). They were further processed to extract higher-level correlation patterns (Figure 5) between single-channel segments (a) and channel-original matches (b). As for AlexNet, the value and saturation components are the most and least correlated to the original, respectively. Its successive checkpoints also tend to be more correlated, which could be due to their coherence emerged through training. ResNet18 shares the same channel ordering but exhibits a stronger correlation level irrespective of checkpoint distances, which can be explained by skip-connections leading to longer distortion retention. The most diverging correlations are observed for checkpoints close to the input (both networks) and fully connected layers (AlexNet). The first pattern can reflect the misalignment in data passed and input layers tuned to specific channels and value ranges. The second case might have arisen due to the suppression of spatial details by adaptive average pooling: this layer decreases the spatial dimensionality and minimizes the overlapping regions, resulting in specific data smoothing, which merges local patterns.
@@ 4 Discussion
This work aimed to design a framework for studying neural networks by exploring the sensitivity of their inner activations to augmentations and auxiliary variables. Sobol indices and Shapley values were used as the main computational tool, which requires a proper sampling scheme for applying augmentations to be employed. Resulting sensitivity maps provide a visualization which captures the individual sensitivity of single units within the checkpoint. Interestingly, the sensitivity to quite simple augmentations is observed up to the deepest layers, which is consistent with the findings from work [20], demonstrating low-level concept identification by deep neurons. Although visual inspection can assist in suggesting the trends, any analysis should be performed with caution due to the approximate nature of the results. We have used complementary ways of analysis through cross-validated LDA, which treats sensitivity maps as features to predict the SA variable, and unit-wise Spearman correlation, which indicates monotonic relationships between variables. Both techniques allowed us to reveal the data patterns at the expense of analysis capability on an individual unit level and showed consistent results. To validate the sensitivity value computations, we performed the guided neural damaging of the corresponding checkpoints. Extracted correlational patterns of matches between the most sensitive classes and top-5 masked predictions demonstrated specificity for a series of augmentations and networks.
The exploration of convolutional neural networks using the constructed framework has established that although a set of augmentations affects all their elements, different layers and neurons respond inhomogeneously to specific SA variables. Such preferences reflect the specialization of network elements that emerged during learning. The specialization changes with the depth of the networks: the closer to the output, the smaller the contribution of augmentations, and the higher the contribution of the class variable. In addition, activation maps reveal on-off patterns similar to those in biological neural networks, which also differ depending on the SA variable and the depth of the layer.
@@
In addition, our findings support the idea of determining single-class sensitivity, which may be potentially utilized to suppress the impact of input distortions on classifying performance. Our approach is based on candidate filtering by co-occurrences of sensible classes in predictions obtained through activation damaging. In general, this is not the only way to gather evidence supporting the correctness of a single-class sensitivity of the whole network by means of the sensitivity of the last classifying layer. Not less important is identifying the classes systematically inhibited by the network with edited activations. It is not obvious how to implement such a method, but it seems to be important as negative feedback and inhibition in biological neural networks.
Analysis of the network segments might be relevant for studying large models with expensive inference, and it is important to recognize the relationships between patterns in isolated and normal operating modes. Through matching the sensitivity measured in both operating modes, it was confirmed that sensitivity to augmentations depends on both input data and preceding transformations. A comparison of correlation patterns among the standalone units also uncovered that skip-connections may facilitate the continuation of the sensitivity patterns. That assumption is also supported by Table 1, according to which the ResNet18 have the smallest percentage decrease in accuracy while applying a single augmentation (the only exception is erasing for VGG11). The method was supplied with a repeating of separate channels of HSV representation, and it may be modified by mapping the values to the expected interval of input values change. Other methods of studying the isolated segments may be proposed, e.g., to generate an activation maximization input and use it as a mask for the corresponding channel, or to generate multiple inputs of specific patterns, like fractal, geometric shapes, textures, noise distributions.
It is worth noting that the proposed sensitivity analysis framework utilizes the classification specifics only for single-class analysis. Auxiliary variables incorporate only the input’s class and partition selection and the order of augmentations to be applied. In other words, they do not rely on the network structure, making it possible to reuse the framework for networks designed to solve other problems.
@@
Results of computational experiments regarding sensitivity values measured at checkpoints of the whole networks and corresponding segments and guided masked predictions are available at the corresponding Zenodo repository:
https://doi.org/10.5281/zenodo.14788913
@@ Acknowledgements
S.M. and I.O. proposed the idea to analyze the sensitivity of neuronal activations using Sobol indices. P.K. suggested studying sensitivity to augmentations and developed the idea into applied research. P.K. and S.M. contributed to the work on the text. I.O. provided administrative support.
@@ Appendix A Experimental setup details
The number of samples to compute the rank-biserial correlation coefficients is calculated as follows: , where and are the number of augmentations with and without inner parameters, - a number of samples for parameterized augmentations, - a number of classes in the dataset, is a number of samples per class. In the case of the sampling scheme 1 (which uses Salteli sampler inside): , where is a number of samples (must be a degree of 2), - number of variables for the -th augmentation set. The number of samples for computing Shapley values is calculated in the following way: , where is a number of permutations, is the similar number of variables for the -th augmentation set, and are the numbers of outer and inner samples.
@@
sharpness (const) | – | – | sharpness factor = 1.5 | |
@@
elliptic local blur | angle (degree) | |||
@@ Appendix B Linear discriminant analysis (LDA) of the convolutional sensitivity maps
Predictions of the SA variables by Sobol indices share the common trends for all networks. The total Sobol indices-based predictions are more diverging, and using the Shapley values results in larger differences in the accuracy dynamics. In the latter case, AlexNet and ResNet18 accuracies evolve similarly on the first augmentation set, and on the second one AlexNet’s and ResNet18’s curves are close for class, partition, contrast variables, VGG11’s and ResNet18’s - for saturation, horizontal flip, rotation, AlexNet’s and VGG11’s - for rotation and saturation. As checkpoints move deeper, the overall accuracies tend to decrease, and there are trend-breaking checkpoints (features.9 for AlexNet, features.10/features.15 for VGG11, layer3 for ResNet18).
Single-variable accuracy patterns are similar for all networks. Erasing, rolling and class variables tend to be the most separable (high accuracy) on the first augmentation set. On the second augmentation set the most discriminative variables are horizontal flip, rotation, class and partition for Shapley values, and elliptic local blur and class for both Sobol indices.
All networks show similar dependencies on confounding with other variables for Shapley values. Class and partition are mutually tangled, and sharping, grayscaling, Gaussian blurring and adjusting brightness, contrast, saturation tend to be predicted instead of permutation. The permutation itself notably draws out the predictions of sharpness and elliptic local blur variables, as well as colour transforms. In general, permutation is confused with almost all augmentations except rolling, where class and partition are more prominent, horizontal flip, which is slightly displaced by hue, and rotation, in which predictions the horizontal flip and hue are occurring. Colour augmentations from the second augmentation set tend to be mutually mispredicted. Predictions of erasing lose their scores due to Gaussian blur, grayscale, sharpness variables; grayscale - by sharpness, Gaussian blur; Gaussian blur - by grayscale, sharpness, partially by erasing.
@@ Appendix C Hierarchical clustering analysis (HCA) of the damaged predictions
By masking activations during the inference and measuring the corresponding accuracies, we obtained prediction patterns, used as a feature space for further correlation analysis. In order to enhance such analysis, we additionally performed the hierarchical clustering procedure, which requires extracting the distances between variables of interest (Figures S7.5(a)–(c)). Computed correlation matrices can be transformed into distance matrices by applying the following transform: , where is a correlation coefficient. HCA has been performed using the hierarchical agglomerative clustering algorithm with average linkage.
@@
In the case of total Sobol indices, the initial distances between single-sample clusters are minimal, and it is difficult to cluster many of them for the first few merges. Thus, the first-order indices and Shapley values are more suitable for further analysis. Pairs with the same masking SA variable tend to group together (especially for ResNet18), which may be explained by a more considerable impact of guided damaging on the inference in comparison with augmentations of input.
@@
Well-matched augmentations and masks of the same variable from the first augmentation set (Gaussian blur and erasing for AlexNet, erasing, grayscale, rolling for VGG11) retain themselves as one-element clusters or ones of a small number of elements until the large-scale merging. There are several exceptions, e.g., Gaussian blur and erasing for ResNet18, potentially indicating its larger robustness to a damage.
For the second augmentation set , all networks tend to early form clusters which include colour adjustment pairs. These maskings tend to be grouped but are entangled in a non-uniform way, so there is no exact subtrees among either networks or sensitivity values.
Appendix D Detailed single-class sensitivity analysis

### v2

@@ 3.3 Cutting off submodules of a network
Corresponding unitwise-averaged correlation matrices are presented in Figures S7.7 (a)–(d) (ImageNet only). They were further processed to extract higher-level correlation patterns (Figure 5) between single-channel segments (a) and channel-original matches (b). As for AlexNet, the value and saturation components are the most and least correlated to the original, respectively. Its successive checkpoints also tend to be more correlated, which could be due to their coherence emerged through training. ResNet18 shares the same channel ordering but exhibits a stronger correlation level irrespective of checkpoint distances, which can be explained by skip-connections leading to longer distortion retention. The most diverging correlations are observed for checkpoints close to the input (both networks) and fully connected layers (AlexNet). The first pattern can reflect the misalignment in data passed and input layers tuned to specific channels and value ranges. The second case might have arisen due to the suppression of spatial details by adaptive average pooling: this layer decreases the spatial dimensionality and minimizes the overlapping regions, resulting in specific data smoothing, which merges local patterns.
@@ 4 Discussion
This work aimed to design a framework for studying neural networks by exploring the sensitivity of their inner activations to simultaneously applied augmentations, enriched with auxiliary variables. Sobol indices and Shapley values were used as the main computational tool, which requires a proper sampling scheme for applying augmentations to be employed. Resulting sensitivity maps provide a visualization which captures the individual sensitivity of single units within the checkpoint. Interestingly, the sensitivity to quite simple augmentations is observed up to the deepest layers, which is consistent with the findings from work [20], demonstrating low-level concept identification by deep neurons. Although visual inspection can assist in suggesting the trends, any analysis should be performed with caution due to the approximate nature of the results. We have used complementary ways of analysis through cross-validated LDA, which treats sensitivity maps as features to predict the SA variable, and unit-wise Spearman correlation, which indicates monotonic relationships between variables. Both techniques allowed us to reveal the data patterns at the expense of analysis capability on an individual unit level and showed consistent results. To validate the sensitivity value computations, we performed the guided neural damaging of the corresponding checkpoints. Extracted correlational patterns of matches between the most sensitive classes and top-5 masked predictions demonstrated specificity for a series of augmentations and networks. Partially the patterns are repeated for fixed architecture on different datasets. Given that we omitted considering fine-tuned networks pretrained on other datasets (transfer learned models), it suggests that sensitivity may be also due to the intrinsic property of the architecture. The resembling patterns within similarly structured AlexNet and VGG11 (ILSVRC) also potentially indicate this.
For a long time the ImageNet was a leading dataset in deep learning, and it continues to be a pretraining basis for various computer vision models until now. In our research, bounded to the ILSVRC (as a subset of ImageNet on training partition and independently sampled on validation part), we thoroughly evaluated developed framework for three classic CNN architectures. The Table 1 implies that all of them still can be improved for a training part and experience an error rate gap between training and validation subsets, thus they are unlikely have surpassed the dataset limitations. Additionally we investigated sensitivity patterns of AlexNet and ResNet18 on the Places365 dataset, similarly to ILSVRC exhibiting the class and imaging conditions diversity, offering a sufficiently high resolution, and featuring accessible pretrained models (excluding VGG11).
Exploration of convolutional neural networks using the constructed framework has established that although a set of augmentations affects all their elements, different layers and neurons respond inhomogeneously to specific SA variables. Such preferences reflect the specialization of network elements that emerged during learning. The specialization changes with the depth of the networks: the closer to the output, the smaller the contribution of augmentations, and the higher the contribution of the class variable. In addition, activation maps reveal on-off patterns similar to those in biological neural networks, which also differ depending on the SA variable and the depth of the layer.
@@
In addition, our findings support the idea of determining single-class sensitivity, which may be potentially utilized to suppress the impact of input distortions on classifying performance. Our approach is based on candidate filtering by co-occurrences of sensitive classes in predictions obtained through activation damaging. In general, this is not the only way to gather evidence supporting the correctness of a single-class sensitivity of the whole network by means of the sensitivity of the last classifying layer. Not less important is identifying the classes systematically inhibited by the network with edited activations. It is not obvious how to implement such a method, but it seems to be important as negative feedback and inhibition in biological neural networks.
Analysis of the network segments might be relevant for studying large models with expensive inference, and it is important to recognize the relationships between patterns in isolated and normal operating modes. Through matching the sensitivity measured in both operating modes, it was confirmed that sensitivity to augmentations depends on both input data and preceding transformations. A comparison of correlation patterns among the standalone units also uncovered that skip-connections may facilitate the continuation of the sensitivity patterns. That assumption is also supported by Table 1, according to which the ResNet18 have the smallest percentage decrease in accuracy while applying a single augmentation (the only exception is erasing for VGG11).
The method was supplied with a repeating of separate channels of HSV representation, and it may be modified by mapping the values to the expected interval of input values change. Other methods of studying the isolated segments may be proposed, e.g., to generate an activation maximization input and use it as a mask for the corresponding channel, or to generate multiple inputs of specific patterns, like fractal, geometric shapes, textures, noise distributions.
It is worth noting that the proposed sensitivity analysis framework utilizes the classification specifics only for single-class analysis. Auxiliary variables incorporate only the input’s class and partition selection and the order of augmentations to be applied. In other words, they do not rely on the network structure, making it possible to reuse the framework for networks designed to solve other problems. Fundamental issue of the variance-based methods is a time-consuming necessity to make a sufficient number of samples. In our implementation, more robust scheme involves saving the sampled values on disc before computing sensivity values, imposing additional requierements on the storage volume.
@@
Another dataset used in our study is Places 365 [54], as it is derived from a large number ( classes) of diverse images of relatively good resolution, and there are available CNN models trained from a scratch on this dataset. Similarly to the ILSVRC, its validation partition is already equalized ( samples per class), and we balanced the number of samples of the training partition to be images per class.
http://places2.csail.mit.edu
Results of computational experiments regarding sensitivity values measured at checkpoints of the whole networks and corresponding segments and guided masked predictions are available at the corresponding Zenodo repositories:
https://doi.org/10.5281/zenodo.18097911
https://doi.org/10.5281/zenodo.18098133
@@ Acknowledgements
S.M. and I.O. proposed the idea to analyze the sensitivity of neuronal activations using Sobol indices. P.K. suggested studying sensitivity to simultaneously applied augmentations, conducted research and developed the idea into the proposed framework. P.K. and S.M. contributed to the work on the text. I.O. provided administrative support.
@@ Appendix A Experimental setup details
The number of samples to compute the rank-biserial correlation coefficients is calculated as follows: , where and are the number of augmentations with and without inner parameters, - a number of samples for parameterized augmentations, - a number of classes in the dataset, is a number of samples per class. In the case of the sampling scheme 1 (which uses Salteli sampler inside): , where is a number of samples (must be a power of 2), - number of variables for the -th augmentation set. The number of samples for computing Shapley values is calculated in the following way: , where is a number of permutations, is the similar number of variables for the -th augmentation set, and are the numbers of outer and inner samples.
@@
| sharpness_const | – | – | sharpness factor = 1.5 | |
@@
| elliptic_local_blur | angle (degree) | |||
@@ Appendix B Linear discriminant analysis (LDA) of the convolutional sensitivity maps
Predictions of the SA variables share several common trends for all networks. Overall accuracies for VGG11 and AlexNet-Places365 tend to form correspondingly higher and lower values, forming the range within which others’accuracy curves are located. Total Sobol indices’ curves tend to be grouped according to the underlying architecture; first-order Sobol ones are less prominent in that sense, and the Shapley values mix AlexNet’s and ResNet18’s curves. As checkpoints progress deeper, the overall accuracies tend to decrease, and there are trend-breaking checkpoints (features.9, AlexNet; features.10/features.15, VGG11; layer3, ResNet18). Shapley values preserve higher overall accuracy values, while they decrease more rapidly in the case of both Sobol indices.
Single-variable accuracy patterns are similar for all networks. Erasing, rolling and class variables tend to be the most separable (high accuracy) on the first augmentation set. On the second augmentation set the most discriminative variables are horizontal flip, rotation, class and partition for Shapley values, and elliptic local blur and class for both Sobol indices. Interestingly, networks tend to be grouped by their architecture and by the dataset in terms of RMS of single-variable accuracies. Even AlexNet trained on Places365 containing many sleeping neurons is closer to either AlexNet trained on ILSVRC than to other networks (Sobol indices) or ResNet18 trained on Places365 dataset (Shapley values). AlexNet-ILSVRC pair is best matching for ResNet18-ILSVRC pair (Shapley) and AlexNet-PLaces365 (total Sobol). ResNet18 prefers its own architecture on both datasets. VGG11 is intermediate - as measured by Shapley values, its patterns appear closer to AlexNet’s on the same dataset, by Sobol indices - to ResNet18’s.
All networks show similar dependencies on confounding with other variables for Shapley values. Class and partition are mutually tangled, and sharping, grayscaling, Gaussian blurring and adjusting brightness, contrast, saturation tend to be predicted instead of permutation. The permutation itself notably draws out the predictions of sharpness and elliptic local blur variables, as well as colour transforms. In general, permutation is confused with almost all augmentations except rolling, where class and partition are more prominent, horizontal flip, which is slightly displaced by hue, and rotation, in which predictions the horizontal flip and hue are occurring. Colour augmentations from the second augmentation set tend to be mutually mispredicted. In addition to permutation, predictions of erasing lose their scores due to Gaussian blur, grayscale, sharpness variables; grayscale - due to sharpness, Gaussian blur; Gaussian blur - grayscale, sharpness, partially erasing.
@@ Appendix C Hierarchical clustering analysis (HCA) of the damaged predictions
By masking activations during the inference and measuring the corresponding accuracies, we obtained prediction patterns, used as a feature space for further correlation analysis. In order to enhance such analysis, we additionally performed the hierarchical clustering procedure, which requires extracting the distances between variables of interest (Figures S7.5.1 (a)–(c), S7.5.2 (a)–(b)). Computed correlation matrices can be transformed into distance matrices by applying the following transform: , where is a correlation coefficient. HCA has been performed using the hierarchical agglomerative clustering algorithm with average linkage.
@@
In the case of total Sobol indices, the initial distances between single-sample clusters are minimal on the ILSVRC dataset, and it is difficult to match many of them for the first few merges. Thus, we restrict the further analysis by the first-order Sobol indices and Shapley values. Pairs with the same masking SA variable tend to group together (especially for ResNet18), which may be explained by a more considerable impact of guided damaging on the inference in comparison with augmentations of input.
@@
Well-matched augmentations and masks of the same variable from the first augmentation set (Gaussian blur, grayscale, and erasing, AlexNet; erasing, grayscale, and rolling, VGG11) retain themselves as one-element clusters or ones of a small number of elements until the large-scale merging. For the second augmentation set , all networks tend to early form clusters which include colour adjustment pairs. These maskings tend to be grouped but are entangled in a non-uniform way, so there is no exact subtrees among either networks or sensitivity values.
Appendix D Selection of the threshold for single-class sensitivity analysis
To isolate the mean Jaccard values for further analysis, we may search for their statistical outliers. The corresponding probabilistic model can be identified as one related to the problem of sampling balls from two identical boxes with unique balls in each. The single Jaccard index corresponds to as a function of , that is, the size of the intersection of two -ball samples, distributed as
In other words, has a hypergeometric distribution with parameters , with equal to the total number of classes, - the number of the most probable classes in a single prediction (top-n). Average value is computed among samples, where is the number of samples per class. Monte-Carlo method of trials provides us with the following estimates: (1) , , : , ; (2) , , : , . Thus, the upper thresholds based on the inter-quantile range, , are: for the ILSVRC dataset, and for Places 365.
Computed thresholds were used to highlight the sufficiently high mean Jaccard indices listed in Tables 5, S7.6.1, S7.6.2. Corresponding outlier detection results are also summarized as Table 4.
| Model | Sens.val. | Augmentation | ||||
| Erasing | Sharp. | Rolling | Grayscale | Gaus.blur | ||
| ILSVRC | ||||||
| AlexNet | Sobol | Table 5 [5/5], S7.6.1 (c) [inv.] | — | S7.6.1 (a) [1/5], (b) [2/5] | Table 5 [3/5] | — |
| Shapley | S7.6.1 (d) [5/5], (e) [2/5] | S7.6.1 (c) [inv.], (d) [5/5], (e) [5/5], (f) [4/5] | S7.6.1 (c) [inv.] | — | S7.6.1 (d) [2/5], (e) [1/5] | |
| VGG11 | Sobol | — | Table 5 [5/5], S7.6.1 (a) [4/5], (b) [5/5], (c) [inv.] | S7.6.1 (a) [3/5], (b) [5/5] | Table 5 [4/5], S7.6.1 (a) [1/5] | Table 5 [1/5], S7.6.1 (a) [1/5] |
| Shapley | S7.6.1 (d) [3/5], (e) [4/5] | — | — | — | — | |
| ResNet18 | Sobol | Table 5 [4/5], S7.6.1 (c) [inv.] | — | S7.6.1 (a) [1/5] | Table 5 [5/5] | — |
| Shapley | — | S7.6.1 (d) [2/5] | — | S7.6.1 (d) [1/5], (f) [2/5] | S7.6.1 (f) [3/5] | |
| Places 365 | ||||||
| AlexNet | Sobol | Table 5 [5/5], S7.6.2 (c) [inv.] | — | Table 5 [1/5], S7.6.2 (a) [1/5] | Table 5 [5/5], S7.6.2 (a) [2/5], (b) [5/5], (c) [inv.] | Table 5 [3/5] |
| Shapley | S7.6.2 (d) [3/5], (e) [5/5] | S7.6.2 (c) [raw], (e) [3/5], (f) [3/5] | S7.6.2 (d) [3/5], (e) [5/5] | S7.6.2 (c) [raw] | S7.6.2 (d) [2/5], (e) [2/5] | |
| ResNet18 | Sobol | Table 5 [5/5], S7.6.2 (c) [inv.] | Table 5 [5/5], S7.6.2 (b) [1/5], (c) [inv.] | Table 5 [1/5], S7.6.2 (a) [2/5], (b) [5/5], (c) [inv.] | — | Table 5 [4/5] |
| Shapley | — | — | S7.6.2 (d) [1/5], (e) [3/5], (f) [5/5] | — | S7.6.2 (c) [inv.], (d) [3/5], (e) [5/5], (f) [5/5] | |
| Dataset | Model | Augmentation | ||||||
| Bright. | Contrast. | Satur. | Hue | Hflip | Rotation | El.loc.blur | ||
| ILSVRC | ||||||||
| AlexNet | Sobol | — | Table 5 [5/5], S7.6.1 (a) [5/5], (b) [4/5] | Table 5 [5/5], S7.6.1 (c) [inv.] | Table 5 [2/5], S7.6.1 (a) [1/5], (b) [5/5], (c) [inv.] | Table 5 [2/5], S7.6.1 (a) [4/5], (b) [4/5], (c) [inv.] | S7.6.1 (a) [1/5], (b) [5/5], (c) [inv.] | Table 5 [3/5], S7.6.1 (a) [4/5] |
| Shapley | — | S7.6.1 (c) [inv.], (d) [4/5], (e) [4/5], (f) [5/5] | S7.6.1 (d) [5/5], (e) [5/5] | — | — | S7.6.1 (f) [4/5] | — | |
| VGG11 | Sobol | S7.6.1 (a) [5/5], (b) [5/5] | Table 5 [5/5], S7.6.1 (a) [2/5], (c) [raw] | Table 5 [5/5], S7.6.1 (a) [5/5], (b) [5/5], (c) [inv.] | S7.6.1 (a) [5/5], (b) [2/5] | S7.6.1 (a) [5/5], (b) [4/5] | Table 5 [2/5], S7.6.1 (a) [3/5], (b) [5/5], (c) [inv.] | S7.6.1 (a) [4/5], (b) [1/5] |
| Shapley | S7.6.1 (c) [inv.], (e) [2/5], (f) [5/5] | S7.6.1 (d) [5/5], (e) [2/5] | S7.6.1 (f) [3/5] | S7.6.1 (d) [5/5] | — | S7.6.1 (c) [raw], (e) [1/5] | S7.6.1 (d) [5/5] | |
| ResNet18 | Sobol | — | Table 5 [5/5], S7.6.1 (a) [1/5], (b) [4/5], (c) [inv.] | Table 5 [5/5], S7.6.1 (b) [1/5] | Table 5 [5/5], S7.6.1 (b) [1/5], (c) [inv.] | — | Table 5 [3/5], S7.6.1 (a) [2/5], (b) [3/5], (c) [inv.] | Table 5 [5/5], S7.6.1 (a) [2/5], (b) [5/5] |
| Shapley | S7.6.1 (c) [inv.], (d) [2/5], (e) [3/5], (f) [5/5] | S7.6.1 (c) [inv.] | S7.6.1 (e) [1/5] | — | S7.6.1 (d) [1/5], (e) [3/5] | S7.6.1 (f) [4/5] | — | |
| Places 365 | ||||||||
| AlexNet | Sobol | Table 5 [5/5], S7.6.2 (b) [1/5], (c) [inv.] | Table 5 [5/5], S7.6.2 (a) [1/5], (b) [1/5], (c) [inv.] | Table 5 [5/5], S7.6.2 (c) [inv.] | Table 5 [5/5], S7.6.2 (c) [inv.] | S7.6.2 (a) [4/5], (b) [4/5], (c) [inv.] | Table 5 [2/5] | — |
| Shapley | S7.6.2 (c) [inv.], (d) [3/5], (e) [2/5], (f) [5/5] | — | S7.6.2 (d) [2/5], (e) [5/5] | S7.6.2 (c) [raw, inv.], (d) [1/5], (e) [3/5] | S7.6.2 (c) [raw], (e) [3/5] | S7.6.2 (c) [inv.], (d) [1/5], (f) [5/5] | S7.6.2 (c) [inv.], (f) [1/5] | |
| ResNet18 | Sobol | S7.6.2 (a) [1/5], (b) [3/5] | S7.6.2 (a) [4/5], (b) [5/5], (c) [inv.] | Table 5 [5/5], S7.6.2 (b) [3/5], (c) [inv.] | Table 5 [5/5], S7.6.2 (c) [inv.] | Table 5 [3/5], S7.6.2 (c) [raw, inv.] | S7.6.2 (a) [4/5], (b) [4/5], (c) [inv.] | S7.6.2 (a) [1/5], (b) [1/5] |
| Shapley | S7.6.2 (c) [inv.], (e) [1/5] | — | — | — | — | S7.6.2 (d) [2/5], (e) [4/5] | — | |
Appendix E Detailed single-class sensitivity analysis